Source sheet: Status_20260831 | Average frequency uses de-duplicated checks from the PanAMA kickoff on 2025-03-04.
| 101 | Izana | operational | RRajagopalan | Sep 10, 2026, 9:56 AM | 44.0 days | |
| 105 | Helsinki | operational | RRajagopalan | Sep 10, 2026, 9:20 AM | 38.1 days | |
| 106 | Innsbruck | operational with issue | MRoca | Sep 10, 2026, 8:30 AM | 48.0 days | |
| 114 | BuenosAires | operational | RRajagopalan | Sep 10, 2026, 8:59 AM | 55.1 days | |
| 115 | Rome-ISAC | operational | HRohringer | Aug 26, 2026, 4:40 PM | 36.7 days | |
| 117 | Rome-SAP | operational | HRohringer | Aug 26, 2026, 4:18 PM | 39.5 days | |
| 120 | Davos | operational | RRajagopalan | Sep 10, 2026, 9:27 AM | 37.7 days | |
| 130 | Lindenberg | out of operation | LHaunold | Sep 10, 2026, 9:17 AM | 56.3 days | no updates from the repair status in berlin |
| 138 | Rome-IIA | operational | HRohringer | Sep 4, 2025, 10:38 AM | 31.0 days | in fieldcampaign in milan |
| 142 | MexicoCity-UNAM | operational with issue | RRajagopalan | Sep 10, 2026, 9:54 AM | 37.7 days | L2 comments: already flagged the data and is caused by temperature so no additional flag needed?
L0 comments: asked MR to put it on hold due to issue until they resolve the fan issue. |
| 149 | Seoul-SNU | operational | RRajagopalan | Sep 10, 2026, 9:03 AM | 38.9 days | L2 comments: Data quality ok but wlshift is going down so might need a calibration soon, needs to revisit in the next QC |
| 162 | Brussels-Uccle | operational with issue | RRajagopalan | Sep 30, 2025, 10:25 AM | 31.6 days | out of operation; BlickM says on the way to SG for repairs and SLC; but we have some files without any data and a new PC in teamviewer; probably arrived in SG; Asked NOps to check. |
| 162 | Utsteinen | operational with issue | LHaunold | Sep 10, 2026, 9:24 AM | 32.3 days | polar night |
| 174 | FairbanksAK | laboratory | LHaunold | Sep 10, 2026, 9:12 AM | 52.8 days | Instrument taken down to be returned after filter wheel began having issues. Full unit being shipped back to SciGlob. |
| 182 | Tel-Aviv | operational | RRajagopalan | Sep 10, 2026, 8:11 AM | 35.6 days | |
| 193 | Tsukuba | hold due to issue | MRoca | Sep 10, 2026, 9:15 AM | 37.0 days | The fiber might be damaged. There is still no response from the PI/LO. |
| 196 | Sapporo | out of operation | MRoca | Sep 10, 2026, 8:47 AM | 45.6 days | Currently the instrument is out of operation. It will be relocated to Bhopal, India |
| 2 | GreenbeltMD | operational | MRoca | Sep 10, 2026, 8:09 AM | 57.8 days | Spec Temp Ok: diurnal rise of temperatures at the limit of the +/-1K threshold, but ok for this instrument since it's hard to stabilize. |
| 209 | Izana | operational | MRoca | Sep 10, 2026, 10:01 AM | 38.2 days | The noticed peaks at noon on fuh5 might be caused by reflections from the surroundings. Still investigating... |
| 235 | Seoul-KU | hold due to issue | RRajagopalan | Sep 10, 2026, 9:11 AM | 35.5 days | Still discussing with SG |
| 240 | Thessaloniki | operational | LHaunold | Sep 10, 2026, 9:21 AM | 33.4 days | some warnings but NOp said only minor warnings |
| 242 | Trollhaugen | operational with issue | RRajagopalan | Sep 10, 2026, 9:42 AM | 35.6 days | |
| 251 | Nainital-ARIES | operational | RRajagopalan | Sep 10, 2026, 8:51 AM | 35.6 days | Pretty bad weather recently so data quality is not great, keep an eye |
| 270 | Warsaw-UW | operational | RRajagopalan | Sep 3, 2026, 3:23 PM | 58.0 days | |
| 280 | Eureka-0PAL | out of operation | RRajagopalan | Sep 10, 2026, 9:13 AM | 42.7 days | HS taken to SciGlob along with spectrometer for repairs on 2026/07/22 |
| 283 | LaPaz | operational | RRajagopalan | Sep 10, 2026, 9:34 AM | 39.6 days | many warnings related to SBHS communication delay (not serious) |
| 289 | DeBilt | operational | MRoca | Sep 10, 2026, 8:54 AM | 37.1 days | |
| 296 | Aosta-SC | operational | HRohringer | Aug 31, 2026, 10:24 AM | 59.3 days | HCHO shows a cycle in the latest period was not here before - will discuss it in operations meeting |
| 309 | Sodankyla | operational | RRajagopalan | Sep 10, 2026, 8:45 AM | 36.0 days | |
| 54 | Seoul | operational with issue | HRohringer | Aug 26, 2026, 4:27 PM | 34.6 days | |
| 68 | WrightwoodCA | operational | RRajagopalan | Sep 10, 2026, 8:43 AM | 41.0 days | L0 comment: Temperature changes a bit during day but within +-1Deg |
| 76 | Dhaka | out of operation | MRoca | Sep 10, 2026, 9:24 AM | 40.6 days | No updates so far
Last NOp note: There was a delay because this shipment was not commercial and required a "diplomatic pouch" to be shipped back to SciGlob. |
| 77 | Singapore-NUS | operational | RRajagopalan | Sep 10, 2026, 9:44 AM | 38.0 days | |
| 80 | BeltsvilleMD | operational | LHaunold | Sep 10, 2026, 9:11 AM | 48.6 days |